### v1

@@ Abstract
The Great Oxidation Event (GOE) was a $200$ Myr transition circa 2.4 billion years ago that converted the Earth’s anoxic atmosphere to one where molecular oxygen (O2) was abundant (volume mixing ratio $10^{-4}$). This significant rise in O2 is thought to have substantially throttled hydrogen (H) escape and the associated water (H2O) loss. Atmospheric estimations from the GOE onward place O2 concentrations ranging between 0.1% to 150% PAL, where PAL is the present atmospheric level of 21% by volume. In this study we use WACCM6, a three-dimensional Earth System Model to simulate Earth’s atmosphere and predict the diffusion-limited escape rate of hydrogen due to varying O2 post-GOE. We find that O2 indirectly acts as a control valve on the amount of hydrogen atoms reaching the homopause in the simulations: less O2 leads to decreased O3 densities, reducing local tropical tropopause temperatures by up to 18 K, which increases H2O freeze-drying and thus reduces the primary source of hydrogen in the considered scenarios. The maximum differences between all simulations in the total H mixing ratio at the homopause and the associated diffusion-limited escape rates are a factor of 3.2 and 4.7, respectively. The prescribed CH4 mixing ratio (0.8 ppmv) sets a minimum diffusion escape rate of $\approx 2\times 10^{10}$ mol H yr-1, effectively a negligible rate when compared to pre-GOE estimates ($\sim 10^{12}-10^{13}$ mol H yr-1). Because the changes in our predicted escape rates are comparatively minor, our numerical predictions support geological evidence that the majority of Earth’s hydrogen escape occurred prior to the GOE. Our work demonstrates that estimations of how the tropical tropopause layer and the associated hydrogen escape rate evolved through Earth’s history requires 3D chemistry-climate models which include a global treatment of water vapour microphysics.
@@
A key component of Earth’s atmospheric and biological history is its water (H2O). The relative amount of H2O that was present during the Earth’s formation versus the quantity delivered by volatile-rich bodies (e.g., comets and asteroids; 2000Icar..148..508D; 2013ApJ...767...54I), as well as the H2O lost to space through hydrogen escape (2001Sci...293..839C; 2008Icar..194...42G), or exchanged into the interior (2015Sci...350..795H; 2017RSPTA.37550393K), has had an important impact on how the Earth’s surface and the atmosphere have developed. Beyond Earth, the solar system’s terrestrial planet atmospheres are desiccated and consequently, Venus and Mars are both apparently desolate environments. It’s possible that Venus was never habitable (2021Natur.598..276T; 2023PNAS..12009751W; 2024NatAs.tmp..289C), but Mars was likely habitable early in its history before its water was isolated to the subsurface and the crust (2010NatGe...3..459D; 2018Sci...361..490O; 2021Sci...372...56S; 2021NatAs...5...63L; 2022NatAs...6.1263S; 2024RemS...16..824Z; wright2024liquid). Venus is now too hot for liquid H2O to persist and probably lost much of its natal H2O through atmospheric hydrogen escape (1969JAtS...26.1191I; 2006P&SS...54.1425K; 2019JGRE..124.2015K; 2023LPICo2891.8035C; Gu_Venus_Photo_2025).
@@
In contrast, Earth has mostly remained temperate throughout its history, with liquid water persisting on the surface for over 4 billion years (2012RvGeo..50.2006F, and references therein), although there were periods of widespread glaciation at either end of the Proterozoic eon (2.4–0.541 Gyr ago; young2013precambrian; 2020PNAS..11713314W). A mostly temperate Earth has been possible despite a fainter Sun in Earth’s past (bahcall2001solar; 2012RvGeo..50.2006F; 2020SSRv..216...90C), in combination with fluctuations in albedo (2020SSRv..216...90C; goldblatt2021earth) and atmospheric composition, in particular the greenhouse gases CO2 (2006PreR..147..148S; 2013ChGeo.362..224S; kanzaki2015estimates) and CH4 (daines2016effect; zhao2018terrestrial; 2019E&PSL.522...48L; fakhraee2019proterozoic). Geochemical proxies suggest a broad range but generally greater past abundance for CO2, whereas only loose constraints have been placed on CH4 (2020SciA....6.1420C, and references therein), with debates regarding whether CH4 during the Proterozoic was higher or lower in abundance still ongoing (see Discussion section).
@@
As all known life requires liquid H2O, the history of Earth’s H2O is inextricably tied to the evolution of Earth’s life. From isotopic evidence and computational modelling, since its formation, the Earth may have lost between 0.13 and 2 times the present ocean volume (2012PNAS..109.4371P; 2018E&PSL.497..149K; 2019GeCoA.244...56Z). Indeed, the early Archean Earth, before 3.2 Gyr ago, may have been fully covered in a global ocean (2021AGUA....200323D). The current hydrogen escape rate is $\approx 4\times 10^{10}$ mol of H $\textrm{yr}^{-1}$, corresponding to roughly 1 metre of global water loss per 1 billion years (2013ChGeo.362...26Z). Given that the Earth’s ocean can be up to 11 km deep (gardner2014so), and its mean depth is $\approx 3.6$ km (charette2010volume), the loss rate is insignificant. This suggests a different past for hydrogen escape than the constant modern rate would imply.
@@
The Archean (4–2.4 Gyr ago) Earth was likely weakly reducing with greater quantities of CH4, CO, and H2 (2001OLEB...31..271K; kharecha2005coupled; 2019GeCoA.262..207K; 2020SciA....6.1420C). Midway through Earth’s history, approximately 2.4 billion years ago, a rise in oxygen known as the Great Oxidation Event (GOE; 2000Sci...289..756F; 2002GeCoA..66.3811H; 2019PNAS..11617207H; 2020PNAS..11713314W; 2021NatGe..14..143G) has been proposed to have significantly reduced hydrogen escape from $\sim 10^{12}-10^{13}$ mol H yr-1 to $\sim 10^{10}$ mol H yr-1 (2001Sci...293..839C; 2013ChGeo.362...26Z). This is because an oxygenated atmosphere destroys reducing species (e.g., H2 and CH4), resulting in lower amounts of such gases which contribute hydrogen atoms to the upper atmosphere.
@@
Whilst hydrogen escape rates may have been much lower in the last $\sim 2.4$ Gyr, to our knowledge, no work has quantitatively estimated escape rates since the GOE occurred using a 3D model. In this work, we build upon a previous study (2022RSOS....911165C) which simulated a range of O2 concentrations that may have been present on the Earth for the last 2.4 Gyr and found that lower O3 columns were predicted compared to previous work (2003AsBio...3..689S; 2015ApJ...806..137R; 2018ApJ...854...19R; 2020ApJ...904...10K; 2022A&A...665A.156K; 2023RSOS...1030056J). More recent work has in part supported the results of lower O3 for a given O2 mixing ratio (2022CliPa..18.2421Y; ji2024correlated; liu2025evolution). O3 and its associated UV heating is responsible for the temperature inversion that creates the Earth’s stratosphere, and this heating affects the temperatures in the TTL. Therefore, the O2 fluctuations following the GOE may have modulated the upward transport of H2O through the TTL and consequently, hydrogen escape too.
@@ 1.1 WACCM6 model
The first simulation we perform commences with initial conditions which are intended to be representative of Earth’s pre-industrial (PI) atmosphere circa 1850, alongside contemporary ocean and land settings. The model’s spatial resolution was set at $1.875^{\circ}$ by $2.5^{\circ}$ (96 latitude points and 144 longitude points), encompassing 70 vertical atmospheric layers spanning from 1000 hPa to $4.5\times 10$-6 hPa (which is up to approximately 150 km in altitude). In total, eight simulations were performed for the Earth system with different mixing ratios of atmospheric O2, where the atmospheric pressure was kept the same by changing the abundance of the background gas, N2. O2 mixing ratios are given in terms of present atmospheric level (PAL), which is 21% by volume. The O2 surface mixing ratios span values between 0.1–150% PAL, which are within estimates for the Earth over the last 2.4 billion years (2014Natur.506..307L; 2019MinDe..54..485L; 2020SciA....6.1420C; 2020PreR..34305722S). These temporal estimates do not always agree, so we remain agnostic and perform simulations covering the whole range. The simulation names are PI, 150% PAL, 50% PAL, 10% PAL, 5% PAL, 1% PAL, 0.5% PAL, and 0.1% PAL. The mixing ratios or fluxes of other gases specified at the surface were kept constant as O2 was altered. This assumption is likely not realistic for the entirety of the last 2.4 Gyr, but it allows us to isolate the role O2 changes make to the chemistry and dynamics of the atmosphere. Water vapor feedback, in terms of evaporation and rainout, is included in WACCM6, but hydrogen escape and water loss are not explicitly simulated. Instead, we estimate these values once quasi-steady state has been reached in the simulation (see section 1.2).
@@
Both atmospheric and oceanic components of the model are interactive, allowing dynamic responses to environmental stimuli like temperature changes. Our simulations incorporated middle atmosphere chemistry as described by 2020JAMES..1201882E, featuring 98 chemical species and 298 reactions, including O3 chemistry. The atmospheric time step, $\Delta t$, was set at 30 minutes. The concentrations of 75 species were computed implicitly, while 22 long-lived species were computed explicitly (sandu1997benchmarking; 2005ama..book.....B) and N2 is considered invariant.
@@
Figure 1: The zonal mean of temperature (in K) is shown for O2 mixing ratios between 150% PAL and 0.1% PAL, where O2 decreases in abundance from the top left panel to the bottom right panel. The colour bar shows hotter temperatures in red and colder temperatures in blue.
@@ 1.2 Diffusion-limited hydrogen escape rate
where $H_{h}$ is the atmospheric scale height at the homopause, $N_{i}$ is the number of hydrogen atoms in each chemical species, $i$, and $b_{\textrm{i}}$ is the binary diffusion parameter for each constituent. The binary diffusion parameters we use are from 1973JAtS...30.1481H and banks1973aeronomy. The binary diffusion coefficient depends on the temperature of the atmosphere at the homopause ($T_{h}$), the masses of the species, and their collision cross-sections. For each species considered here, the binary diffusion coefficients are $b_{\textrm{H}}=6.5\times 10^{17}T_{h}^{0.7}$, $b_{\textrm{H}_{2}}=2.67\times 10^{17}T_{h}^{0.75}$, $b_{\textrm{H}_{2}\textrm{O}}=0.137\times 10^{17}T_{h}^{1.072}$, $b_{\textrm{CH}_{4}}=0.756\times 10^{17}T_{h}^{0.747}$. The calculations are made for the whole atmosphere (latitude and longitude at the homopause) before calculating a longitudinally-averaged and latitudinally-weighted mean.
@@ 2.1 Water vapour through the cold trap
Figure 2: The mean tropical (defined at latitudes $\pm 24^{\circ}$ from the equator) O3 mixing ratio between 50–450 hPa is plotted on the left vertical axis in teal circles against the atmospheric mixing ratio of O2 at the surface in terms in terms of the present atmospheric level (PAL), which is 21% by volume. The total global ozone (O3) column density in Dobson Units (DU, where 1 DU $2.69\times 10^{20}$ molecules m-2) is also plotted in purple squares on the right vertical axis against the atmospheric mixing ratio of O2.
@@
In every simulation, the global mean mixing ratio of each species at the surface for H2O, CH4, H2, and H is $\sim 10,000$ ppmv, $\approx 0.8$ ppmv, $\approx 0.5$ ppmv and $10^{-8}$ ppmv respectively. As temperature and pressure decline with altitude in the troposphere, the atmosphere cannot hold as much H2O vapor. This is because the drop in pressure causes the air to cool (adiabatic cooling), and the saturation vapor pressure of water, which determines the air’s capacity for water vapor, is strongly dependent on temperature. Fig. 1 presents the zonal mean (averaged over longitude) of the atmospheric temperature between the surface and 10 hPa ($\sim 30$ km in altitude) for all the WACCM6 simulations. The temperature structure of the atmosphere is set by the strength and spectral shape of the incoming radiation, the moist adiabat in the troposphere, atmospheric composition (providing heating and cooling in different layers), clouds and ice, and atmospheric dynamics. Between 150% PAL and 50% PAL, the temperature structure at pressures greater than (altitudes below) 10 hPa is similar. At 10% and 5% PAL, the temperature increases in the middle atmosphere, and this is noticeable in the tropics between 200–30 hPa and at northern ($45-90^{\circ}$) latitudes between 200–100 hPa. At $\leq 1$% PAL of O2, the stratosphere begins to become significantly cooler, with the 10 hPa pressure level $\sim 60$ K cooler in the 0.1% PAL simulation compared to the 150% PAL simulation. A globally averaged stratospheric temperature inversion of $\approx 2$ K and $\approx 1$ K exists in the 1% PAL and 0.1% PAL simulations, respectively, which is significantly smaller than the $\approx 60$ K temperature inversion in the PI atmosphere.
Fig. 2 shows the mixing ratio of atmospheric O2 plotted against the O3 column density, as well as the mean O3 mixing ratio between 50 and 450 hPa in the tropical region. Whilst the global O3 column density increases with increasing O2 mixing ratio, this is not the case for the O3 mixing ratio around the TTL region: there is instead a peak at $\approx$ 10% PAL. O3 in this region absorbs UV light and heats the TTL, which is the cause of the hotter temperatures in the higher oxygenated scenarios in Fig. 1 compared to scenarios with <1<1% PAL. This is also why there is an asymmetry in temperatures because of hemispherical asymmetries in the O3 distribution (there is more O3 in the northern hemisphere) as a result of the Brewer-Dobson circulation (1926RSPSA.110..660D; 1949QJRMS..75..351B; 2014RvGeo..52..157B).
@@
In Section The Oxygen Valve on Hydrogen Escape Since the Great Oxidation Event, we described how H2O moves through the TTL to get into the stratosphere on modern Earth. The cold trap mechanism is indirectly sensitive to O2 concentrations resulting from the aforementioned difference in O3 and hence UV heating. Due to the difference in time-averaged zonal mean TTL temperatures between the PI and 0.1% PAL simulations of up to $\approx 18$ K, more H2O is frozen out in the form of water clouds, ice, and ice clouds before crossing $\approx 100$ hPa in the 0.1% PAL case compared to the standard PI case (2012JGRD..117.4211W). Stratospheric O3 and H2O are greenhouse gases, and with less H2O reaching the stratosphere as a consequence of reduced O3, this decreases tropospheric heating (2023CliPa..19.1201D), causing a positive feedback effect, albeit a small one: there is some cooling at the surface, but this is limited to <3<3 K in terms of global averaged surface temperatures.
@@
Figure 3: The top left, top right, and bottom left panels show the high cloud fraction as a function of longitude and latitude for the PI, 1% PAL, and 0.1% PAL scenarios, respectively. Bottom right: The high cloud fraction (high clouds are defined as clouds at pressures <400<400 hPa) is averaged over longitude and shown as a function of latitude in the top panel for all the WACCM6 simulations.
Figure 4: The atmospheric ice content (blue-white shading) and the amount of atmospheric H2O (coloured contours) is shown on the latitude-longitude grid of the Earth. Both are given in terms of ppmv. The PI simulation (left) and the 0.1% PAL simulation are shown (right). Both scenarios are shown for a pressure level of 142 hPa, which corresponds to an altitude of $\approx 14$ km near the base of the TTL.

### v2

@@ Abstract
The Great Oxidation Event (GOE) was a $200$ Myr transition circa 2.4 billion years ago that converted the Earth’s anoxic atmosphere to one where molecular oxygen (O2) was abundant (volume mixing ratio $10^{-4}$). This significant rise in O2 is thought to have substantially throttled hydrogen (H) escape and the associated water (H2O) loss. Atmospheric estimations from the GOE onward place O2 concentrations ranging between 0.1% to 150% PAL, where PAL is the present atmospheric level of 21% by volume. In this study we use WACCM6, a three-dimensional Earth System Model to simulate Earth’s atmosphere and predict the diffusion-limited escape rate of hydrogen due to varying O2 post-GOE. We find that O2 indirectly acts as a control valve on the amount of hydrogen atoms reaching the homopause in the simulations: less O2 leads to decreased O3 densities that reduce local tropical tropopause temperatures by up to 17 K, which increases H2O freeze-drying and thus reduces the primary source of hydrogen in the considered scenarios. The maximum differences between all simulations in the total H mixing ratio at the homopause and the associated diffusion-limited escape rates are a factor of 3.2 and 4.7, respectively. The prescribed CH4 mixing ratio (0.8 ppmv) sets a minimum diffusion escape rate of $\approx 2\times 10^{10}$ mol H yr-1, effectively a negligible rate when compared to pre-GOE estimates ($\sim 10^{12}-10^{13}$ mol H yr-1). Because the changes in our predicted escape rates are comparatively minor, our numerical predictions support geological evidence that the majority of Earth’s hydrogen escape occurred prior to the GOE. Our work demonstrates that estimations of how the tropical tropopause layer and the associated hydrogen escape rate evolved through Earth’s history requires 3D chemistry-climate models which include a global treatment of water vapour microphysics.
@@
A key component of Earth’s atmospheric and biological history is its water (H2O). The relative amount of H2O that was present during the Earth’s formation versus the quantity delivered by volatile-rich bodies (e.g., comets and asteroids; 2000Icar..148..508D; 2013ApJ...767...54I), as well as the H2O lost to space through hydrogen escape (2001Sci...293..839C; 2008Icar..194...42G), or exchanged into the interior (2015Sci...350..795H; 2017RSPTA.37550393K), has had an important impact on how the Earth’s surface and the atmosphere have developed. Beyond Earth, the solar system’s terrestrial planet atmospheres are desiccated and consequently, Venus and Mars are both apparently desolate environments. It’s possible that Venus was never habitable (2021Natur.598..276T; 2023PNAS..12009751W; 2024NatAs.tmp..289C), but Mars was likely habitable early in its history before its water was isolated to the subsurface and the crust (2010NatGe...3..459D; 2018Sci...361..490O; 2021Sci...372...56S; 2021NatAs...5...63L; 2022NatAs...6.1263S; 2024RemS...16..824Z; wright2024liquid). Venus is now too hot for liquid H2O to persist and probably lost much of its natal H2O through atmospheric hydrogen escape (1969JAtS...26.1191I; 2006P&SS...54.1425K; 2019JGRE..124.2015K; 2024AGUFMSM13E..08C; Gu_Venus_Photo_2025).
@@
In contrast, Earth has mostly remained temperate throughout its history, with liquid water persisting on the surface for over 4 billion years (2012RvGeo..50.2006F, and references therein), although there were periods of widespread glaciation at either end of the Proterozoic eon (2.4 – 0.541 Gyr ago; young2013precambrian; 2020PNAS..11713314W). A mostly temperate Earth has been possible despite a fainter Sun in Earth’s past (bahcall2001solar; 2012RvGeo..50.2006F; 2020SSRv..216...90C), in combination with fluctuations in albedo (2020SSRv..216...90C; goldblatt2021earth) and atmospheric composition, in particular the greenhouse gases CO2 (2006PreR..147..148S; 2013ChGeo.362..224S; kanzaki2015estimates) and CH4 (daines2016effect; zhao2018terrestrial; 2019E&PSL.522...48L; fakhraee2019proterozoic). Geochemical proxies suggest a broad range but generally greater past abundance for CO2, whereas only loose constraints have been placed on CH4 (2020SciA....6.1420C, and references therein), with debates regarding whether CH4 during the Proterozoic was higher or lower in abundance than present day still ongoing (see Discussion section).
@@
As all known life requires liquid H2O, the history of Earth’s H2O is inextricably tied to the evolution of Earth’s life. From isotopic evidence and computational modelling, since its formation, the Earth may have lost between 0.13 and 2 times the present ocean volume (2012PNAS..109.4371P; 2018E&PSL.497..149K; 2019GeCoA.244...56Z). Indeed, the early Archean Earth (4 – 2.4 Gyr ago), before 3.2 Gyr ago, may have been fully covered in a global ocean (2021AGUA....200323D). The current hydrogen escape rate is $\approx 4\times 10^{10}$ mol of H $\textrm{yr}^{-1}$, corresponding to roughly 1 metre of global water loss per 1 billion years (2013ChGeo.362...26Z). Given that the Earth’s ocean can be up to 11 km deep (gardner2014so), and its mean depth is $\approx 3.6$ km (charette2010volume), the loss rate is insignificant. This suggests a different past for hydrogen escape than the constant modern rate would imply.
@@
The atmosphere of the Archean Earth was likely weakly reducing with greater quantities of CH4, CO, and H2 (2001OLEB...31..271K; kharecha2005coupled; 2019GeCoA.262..207K; 2020SciA....6.1420C). Midway through Earth’s history, approximately 2.4 billion years ago, a rise in oxygen known as the Great Oxidation Event (GOE; 2000Sci...289..756F; 2002GeCoA..66.3811H; 2019PNAS..11617207H; 2020PNAS..11713314W; 2021NatGe..14..143G) has been proposed to have significantly reduced hydrogen escape from $\sim 10^{12}-10^{13}$ mol H yr-1 to $\sim 10^{10}$ mol H yr-1 (2001Sci...293..839C; 2013ChGeo.362...26Z). This is because an oxygenated atmosphere destroys reducing species (e.g., H2 and CH4), resulting in lower amounts of such gases which contribute hydrogen atoms to the upper atmosphere.
@@
Whilst hydrogen escape rates may have been much lower in the last $\sim 2.4$ Gyr, to our knowledge, no work has quantitatively estimated escape rates since the GOE occurred using a 3D model. In this work, we build upon a previous study (2022RSOS....911165C) which simulated a range of O2 concentrations that may have been present on the Earth for the last 2.4 Gyr and found that lower O3 columns were predicted compared to previous work (2003ARA&A..41..429K; 2003AsBio...3..689S; 2015ApJ...806..137R; 2018ApJ...854...19R; 2020ApJ...904...10K; 2023RSOS...1030056J). More recent work has in part supported the results of lower O3 for a given O2 mixing ratio (2022CliPa..18.2421Y; ji2024correlated; liu2025evolution). O3 and its associated UV heating is responsible for the temperature inversion that creates the Earth’s stratosphere, and this heating affects the temperatures in the TTL. Therefore, the O2 fluctuations following the GOE may have modulated the upward transport of H2O through the TTL and consequently, hydrogen escape too.
@@ 1.1 WACCM6 model
The first simulation we perform commences with initial conditions which are intended to be representative of Earth’s pre-industrial (PI) atmosphere circa 1850, alongside contemporary ocean and land settings, and the current orbital setup with a $23.4^{\circ}$ obliquity and 24-hour rotation rate. Deep time paleoclimate studies would often change the continents and we accept that this is a limitation with the WACCM6 model where continents cannot currently be modified. The circulation pattern and climate would have been different in the past due to various factors, which include the younger Sun, different greenhouse gas concentration, and the formation and break up of supercontinents such as Pangea and Columbia/Nuna (Rogers_2002_Columbia; 2012RvGeo..50.2006F; 2017Geo....45..231F; 2020SciA....6.1420C). The model’s spatial resolution for the atmosphere was set at $1.875^{\circ}$ latitude by $2.5^{\circ}$ longitude (yielding a $96\times 144$ horizontal grid). This configuration utilizes 70 atmospheric layers in vertical pressure coordinates extending from the surface (approximately 1000 hPa) to the thermosphere ($4.5\times 10^{-6}$ hPa, or $\sim 150$ km). The land component (CLM5.0) operates on the same $1.9^{\circ}\times 2.5^{\circ}$ horizontal grid as the atmosphere to maintain consistency in surface flux exchanges. The simulation employs an interactive ocean via the POP2 model, which operates on a $1^{\circ}$ displaced-pole grid. This ocean grid features $320\times 384$ points, with enhanced latitudinal resolution near the equator to better resolve tropical dynamics. This same grid is utilized by the Sea-Ice (CICE5) and Wave (WaveWatch3) components. Additionally, freshwater discharge is included in the MOSART river transport model on a $0.5^{\circ}$ global grid.
@@
Both atmospheric and oceanic components of the model are interactive, allowing dynamic responses to environmental stimuli like temperature changes. Our simulations incorporated middle atmosphere chemistry as described by 2020JAMES..1201882E, featuring 98 chemical species and 298 reactions, including O3 chemistry. The atmospheric time step, $\Delta t$, was set at 30 minutes. The concentrations of 75 species were computed using an implicit numerical scheme, while 22 long-lived species were computed using an explicit numerical scheme (sandu1997benchmarking; 2005ama..book.....B). N2 is considered invariant. Absorption of light by CO2 and H2O in the Schumann-Runge (S-R bands) is not included in these simulations, and neither is scattering at wavelengths <200<200 nm. For more details on the inclusion of these processes, we refer the reader to ji2023comparison and ji2024correlated, as well as Section 3.2 where we discuss how it might impact our results.
In total, eight simulations were performed for the Earth system with different mixing ratios of atmospheric O2, where the atmospheric pressure was kept the same by changing the abundance of the background gas, N2. O2 mixing ratios are given in terms of present atmospheric level (PAL), which is 21% by volume. The O2 surface mixing ratios span values between 0.1 – 150% PAL, which are within estimates for the Earth over the last 2.4 billion years (2014Natur.506..307L; 2019MinDe..54..485L; 2020SciA....6.1420C; 2020PreR..34305722S). These temporal estimates do not always agree, so we remain agnostic and perform simulations covering the whole range. The simulation names are PI, 150% PAL, 50% PAL, 10% PAL, 5% PAL, 1% PAL, 0.5% PAL, and 0.1% PAL. CO2, CH4, N2O, and H2, were held fixed at values of 284 ppmv, 0.8 ppmv, 273 ppbv, and 0.5 ppmv, respectively. The mixing ratios or fluxes of gases specified at the surface were kept constant as O2 was altered. This assumption is likely not realistic for the entirety of the last 2.4 Gyr, but it allows us to isolate the role O2 changes make to the chemistry and dynamics of the atmosphere. Water vapor feedback, in terms of evaporation and rainout, is included in WACCM6, but hydrogen escape and water loss are not explicitly simulated. Instead, we estimate these values once quasi-steady state has been reached in the simulation (see section 1.2).
@@ 1.2 Diffusion-limited hydrogen escape rate
where $H_{h}$ is the atmospheric scale height at the homopause, $N_{i}$ is the number of hydrogen atoms in each chemical species, $i$, and $b_{\textrm{i}}$ is the binary diffusion parameter for each constituent. The binary diffusion coefficient depends on the temperature of the atmosphere at the homopause ($T_{h}$), the masses of the species, and their collision cross-sections. For each species considered here, the binary diffusion coefficients are $b_{\textrm{H}}=6.5\times 10^{17}T_{h}^{0.7}$, $b_{\textrm{H}_{2}}=2.67\times 10^{17}T_{h}^{0.75}$, $b_{\textrm{H}_{2}\textrm{O}}=0.137\times 10^{17}T_{h}^{1.072}$, $b_{\textrm{CH}_{4}}=0.734\times 10^{17}T_{h}^{0.750}$. These binary diffusion parameters for H, H2, and H2O are sourced from 1973JAtS...30.1481H and CH4 is sourced from banks1973aeronomy. The calculations are made for the whole atmosphere (latitude and longitude at the homopause) before calculating a longitudinally-averaged and latitudinally-weighted mean.
@@ 2 Results
Figure 1: The zonal mean of atmospheric temperature (in K) is shown for simulations with O2 mixing ratios between 150% PAL and 0.1% PAL, where O2 decreases in abundance from the top left panel to the bottom right panel. The colour bar shows hotter temperatures in red and colder temperatures in blue. The TTL region is $\pm 23.4^{\circ}$ from the equator and between 70–150 hPa in pressure.
@@ 2.1 Water vapour through the cold trap
Figure 2: Left: The mean tropical (defined at latitudes $\pm 24^{\circ}$ from the equator) O3 mixing ratio between 70 – 150 hPa is plotted on the left vertical axis in teal circles against the atmospheric mixing ratio of O2 at the surface in terms in terms of the present atmospheric level (PAL), which is 21% by volume. The total global ozone (O3) column density in Dobson Units (DU, where 1 DU $2.69\times 10^{20}$ molecules m-2) is also plotted in purple squares on the right vertical axis against the atmospheric mixing ratio of O2. Right: The O3 number density is shown for all eight simulations between 10 – 200 hPa on a linear scale (note the number density does not go to zero at pressures greater than 200 hPa, see 2022RSOS....911165C).
@@
In every simulation, the global mean mixing ratio of each species at the surface for H2O, H, CH4, and H2, is $\sim 10,000$ ppmv, $10^{-8}$ ppmv, $\approx 0.8$ ppmv, and $\approx 0.5$ ppmv respectively, with the latter two molecules specified with fixed mixing ratios at the surface. As temperature and pressure decline with altitude in the troposphere, the atmosphere cannot hold as much H2O vapor. This is because the drop in pressure causes the air to cool (adiabatic cooling), and the saturation vapor pressure of water, which determines the air’s capacity for water vapor, is strongly dependent on temperature. Fig. 1 presents the zonal mean (averaged over longitude) of the atmospheric temperature between the surface and 10 hPa ($\sim 30$ km in altitude) for all the WACCM6 simulations. The temperature structure of the atmosphere is set by the strength and spectral shape of the incoming radiation, the moist adiabat in the troposphere, atmospheric composition (providing heating and cooling in different layers), clouds and ice, and atmospheric dynamics. Between 150% PAL and 50% PAL, the temperature structure at pressures greater than 10 hPa (or altitudes below 10 hPa) is similar. At 10% and 5% PAL, the temperature increases in the middle atmosphere, and this is noticeable in the tropics between 200 – 30 hPa and at northern ($45-90^{\circ}$) latitudes between 200 – 100 hPa. At $\leq 1$% PAL of O2, the stratosphere begins to become significantly cooler, with the 10 hPa pressure level $\sim 60$ K cooler in the 0.1% PAL simulation compared to the 150% PAL simulation. A globally averaged stratospheric temperature inversion of $\approx 2$ K and $\approx 1$ K exists in the 1% PAL and 0.1% PAL simulations, respectively, which is significantly smaller than the $\approx 60$ K temperature inversion in the PI atmosphere.
Fig. 2 shows the mixing ratio of atmospheric O2 plotted against the O3 column density, as well as the mean O3 mixing ratio between 70 and 150 hPa in the tropical region, accounting for the expanse of the TTL. The figure also shows the O3 number density in the tropics between 10 and 200 hPa. Whilst the global O3 column density increases with increasing O2 mixing ratio, this is not the case for the O3 mixing ratio around the TTL region: there is instead a peak at $\approx 5-10$% PAL. The O3 number density also mirrors this relationship. When O2 is decreased, O3 accumulates lower down in the atmosphere because the wavelengths that photolyse O2 to produce “odd oxygen” (O + O3) are able to penetrate deeper. Consequently, for O2 concentrations of $\leq 10$% PAL, O3 maximizes in this region near the TTL. The TTL temperatures are a balance between shortwave heating and the emission and absorption of longwave radiation. Whilst this causes higher temperatures than the PI atmosphere for 5 – 10% PAL, simulations with $\leq 1$% PAL have resultant lower temperatures. This effect is a local TTL phenomenon; globally, the mean tropopause temperatures decrease with decreasing O2.
These results are in line with previous O3 depletion experiments (albeit of a much smaller magnitude) which found that reduced O3 concentrations cooled the TTL and led to less stratospheric water vapour (Forster_2007_O3; Xie_2008_O3_tape; Lu_2021_O3; Zhou_2022_Ozone). The cooling was attributed to radiative processes. Indeed, Forster_2007_O3 showed that less O3 in the tropical lower stratosphere caused cooling that extended into the upper tropical troposphere due to reduced downward longwave radiation. Similarly, here we find that lower middle stratospheric densities of O3 reduces the propagation of longwave radiation downwards. This mechanism is the cause of the hotter temperatures in the higher oxygenated scenarios in Fig. 1 compared to scenarios with $\leq 1$% PAL. This is also why there is an asymmetry in temperatures because of hemispherical asymmetries in the O3 distribution (there is more O3 in the northern hemisphere) as a result of the Brewer-Dobson circulation (1926RSPSA.110..660D; 1949QJRMS..75..351B; 2014RvGeo..52..157B).
@@
In Section The Oxygen Valve on Hydrogen Escape Since the Great Oxidation Event, we described how H2O moves through the TTL to get into the stratosphere on modern Earth. The cold trap mechanism is indirectly sensitive to O2 concentrations resulting from the aforementioned difference in O3 and hence UV heating. Due to the difference in time-averaged zonal mean TTL temperatures between the PI and 0.1% PAL simulations of up to $\approx 17$ K, more H2O is frozen out in the form of water clouds, ice, and ice clouds (e.g., cirrus clouds; 2012JGRD..117.4211W) before crossing $\approx 100$ hPa in the 0.1% PAL case compared to the standard PI case. Stratospheric O3 and H2O are greenhouse gases, and with less H2O reaching the stratosphere as a consequence of reduced O3, this decreases tropospheric heating (2023CliPa..19.1201D), causing a positive feedback effect, albeit a small one: there is some cooling at the surface, but this is limited to <3<3 K in terms of global averaged surface temperatures.
@@
Figure 3: The top left, top right, and bottom left panels show the high cloud fraction as a function of longitude and latitude for the PI, 1% PAL, and 0.1% PAL scenarios, respectively. Bottom right: The high cloud fraction (high clouds are defined as clouds at pressures <400<400 hPa) is averaged over longitude and shown as a function of latitude for all the WACCM6 simulations.
Figure 4: The atmospheric ice content (blue-white shading) and the amount of atmospheric H2O (magenta-orange-yellow contours) is shown on the latitude-longitude grid of the Earth. Both are given in terms of ppmv. The PI simulation (left) and the 0.1% PAL simulation are shown (right). Both scenarios are shown for a pressure level of 142 hPa, which corresponds to an altitude of $\approx 14$ km near the base of the TTL.